OpenStudy (anonymous):

A water sprinkler has a range of 5 meters as shown. The jet of water from the sprinkler sweeps out at an angle of 85° as the nozzle turns. What is the area watered by the sprinkler? Use the value π = 3.14, and round your answer to one decimal place. a. 15.6 Square Meters b. 17.8 Square Meters c. 18.5 Square Meters d. 19.2 Square Meters

OpenStudy (anonymous):

try PI * R^2 so 3.14 * 5^2 then subtract 85 degrees from your answer

OpenStudy (anonymous):

Why would I subtract 85 degrees?

OpenStudy (anonymous):

well im not sure yet because im debating against myself if you should divide or subtract sorry. :) so maybe try both

OpenStudy (anonymous):

3.14 * 5^2 = 78.5 78.5 - 85 = -6.5 So that would be wrong 85 / 75 = 1.083 Something is missing I think, I'm just not sure what exactly

OpenStudy (anonymous):

I have to log off schools dismissed but thank you anyway